### v1

# Fast Inference of Removal-Based Node Influence

Abstract.
Graph neural networks (GNNs) are widely utilized to capture the information spreading patterns in graphs. While remarkable performance has been achieved, there is a new trending topic of evaluating node influence. We propose a new method of evaluating node influence, which measures the prediction change of a trained GNN model caused by removing a node. A real-world application is, “In the task of predicting Twitter accounts’ polarity, had a particular account been removed, how would others’ polarity change?”. We use the GNN as a surrogate model whose prediction could simulate the change of nodes or edges caused by node removal. To obtain the influence for every node, a straightforward way is to alternately remove every node and apply the trained GNN on the modified graph. It is reliable but time-consuming, so we need an efficient method. The related lines of work, such as graph adversarial attack and counterfactual explanation, cannot directly satisfy our needs, since they do not focus on the global influence score for every node. We propose an efficient and intuitive method, NOde-Removal-based fAst GNN inference (NORA), which uses the gradient to approximate the node-removal influence. It only costs one forward propagation and one backpropagation to approximate the influence score for all nodes. Extensive experiments on six datasets and six GNN models verify the effectiveness of NORA. Our code is available at https://github.com/weikai-li/NORA.git.
Graph Neural Network, Web Mining, Node Influence Evaluation
copyright: acmcopyrightjournalyear: 2024copyright: rightsretainedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589334.3645389isbn: 979-8-4007-0171-9/24/05submissionid: 434ccs: Information systems Web miningccs: Computing methodologies Neural networks
1. Introduction
Measuring node influence in a graph and identifying influential nodes are important to various applications, such as advertising (Domingos and Richardson, 2001; Nguyen et al., 2017; Leskovec et al., 2007a), online news dissemination (Leskovec et al., 2007b; Gruhl et al., 2004), finding bottlenecks in an infrastructure network to improve robustness (Li and Liu, 2021; Crucitti et al., 2004), vaccination on prioritized groups of people to break down virus spreading (Bellingeri et al., 2020; Holme, 2004; Wang et al., 2015), etc. This topic has attracted many studies. “Influence maximization” problem aims to identify influential nodes whose triggered influence spreading range is maximized (Li et al., 2022; Lin et al., 2020; Huang et al., 2019; Jalayer et al., 2018; Sun et al., 2022; Zhang et al., 2022b; Tian et al., 2020; Ling et al., 2023). “Network dismantling” problem studies the influence of node removal on network connectivity(Nguyen et al., 2019; Mburano et al., 2022; Lou et al., 2022; Osat et al., 2022; Yan et al., 2022; Zhang and Wang, 2022).
Figure 1. An example of the task-specific influence of node removal in social networks. Red versus Blue represents two different opinions, and color shades represent the degree of opinion. When the top blue node is removed, the two pink nodes hear less voice from the blue nodes and become red. The two left nodes no longer follow the middle node, and the left white node becomes blue. These are the influence of removing the top blue node.
They define the node influence based on the graph structure (e.g., connectivity) or a given propagation model (e.g., susceptible-infected-removed model). However, these definitions are not flexible enough to capture the node influence from different aspects. For example, we might want to identify the biggest political influencers on Twitter. In this case, we want to calculate the influence score of a Twitter account based on how much it would affect other users’ political polarity had it been removed. In another scenario, we might want to identify the biggest fashion influencers on Twitter, and we want to calculate the node influence based on how much it would affect other users’ fashion categories had it been removed. Instead of adopting any fixed node influence definition, we thus focus on task-specific node influence score calculation based on node removal. Figure 1 provides an example of the task-specific node influence.
Graph neural networks (GNNs) are among the most powerful graph learning tools. We use GNNs as a surrogate to capture the underlying mechanism of how the graph structure affects node behaviors. In the ideal case, we should train a new GNN based on the graph where the target node is removed and other node/edge labels could also be different. Unfortunately, this graph only exists in a parallel world and is not available for training GNNs. We have two options to solve this issue. First, we can assume the model does not change significantly, so we can use the model trained on the original graph. Second, we can assume the labels of other nodes/edges do not change, so we can re-train the GNN on the new graph where the target node is removed and the labels do not change. We choose the first option, because the underlying patterns of message spreading learned by the GNN should be relatively stable. After removing a node, we use the new predictions of the trained GNN on the modified graph to simulate the new labels in the parallel world. We calculate the influence of node removal as the total variation distance between the original predictions and new predictions, as illustrated in Figure 2.
Figure 2. Our schema of calculating node influence. The GNN model is trained on the original graph. We remove a node and apply the trained GNN to the modified graph. We calculate the total variation distance between the original predictions and new predictions as the influence of node removal.
Our target is to calculate the influence score for every node. The most straightforward way is to alternately remove each node from the original graph and use the trained GNN to do prediction. However, it is very time consuming, so we demand an efficient method. Evaluating the change of GNN predictions caused by the change of input has been studied in graph adversarial attack and graph counterfactual explanation. Graph adversarial attack aims to maximally undermine GNN performance or change GNN prediction by perturbations to the input graph, which mainly include modifying node features (Zügner et al., 2018, 2020; Ma et al., 2021), injecting nodes (Sun et al., 2020; Wang et al., 2020; Zou et al., 2021; Chen et al., 2022; Dai et al., 2020; Ju et al., 2022a; Tao et al., 2021), or edge perturbation (Zügner and Günnemann, 2019; Xu et al., 2019; Zhang et al., 2022a; Wang et al., 2023). However, to the best of our knowledge, none of the adversarial attack methods utilizes node removal, since it is not common in the attack scenario.
Graph counterfactual explanation aims to explain the GNN’s prediction of a target node/edge/graph by finding the minimum perturbation on the input graph that can change the prediction of the target (Stepin et al., 2021). Some methods utilize node removal (Tan et al., 2022; Wu et al., 2021; Rong et al., 2023; Ma et al., 2022; Yuan et al., 2020, 2021; Huang et al., 2023), but they can not directly satisfy our needs. First, our problem setting is different. We evaluate the influence of removing one node on other nodes/edges, while the explanation methods focus on the influence of removing several nodes on one target node/edge/graph. Second, the existing strategies can not easily scale up to handle large graphs when the goal is to predict the influence score for every node. We briefly summarize the difference in problem settings between related lines of research and this paper in Table 1. There are three important aspects in our problem setting: task-specific, influence of node removal, and global influence (influence on all nodes/edges). As shown in the table, the existing lines of work generally do not simultaneously have these three aspects.
Table 1. Differences in problem settings.
Task-specific Removal Global influence
Influence maximization
Network dismantling
Adversarial attack
Counterfactual explanation
This paper
To efficiently calculate the node influence score, we propose an intuitive, effective, and efficient method, NOde-Removal-based fAst GNN Inference (NORA). We use the first-order derivatives to approximate the influence. It is model-agnostic and can be easily adapted to any common GNN model based on the message-passing framework. It only needs one forward propagation and one backpropagation to approximate the removal influence for all nodes. It takes up to 41 hours to generate the real node influence in our experiments, while NORA uses less than a minute if we do not include the time of generating the validation-set labels. Although simple and intuitive, NORA works well in our experiments. We modify and adapt two approaches in graph counterfactual explanation as baselines, and NORA outperforms them in the experiments. To sum up, this paper makes the following contributions:
  • New Problem. We propose a novel perspective of evaluating task-specific node influence based on node removal by GNN.
  • Methodology. We propose an efficient and effective algorithm, NORA, to approximate the node influence for all nodes.
  • Effectiveness. Experimental results on six datasets and six GNN models demonstrate that NORA outperforms the baselines.
2. Related Work
2.1. Graph Adversarial Attack
Graph adversarial attack aims to maximally undermine GNN performance or change GNN predictions by imposing a small perturbation. Zügner et al. (Zügner et al., 2018, 2020) started the race of graph adversarial attacks. Pioneering works mainly focused on modifying node features (Zügner et al., 2018, 2020; Ma et al., 2021) and perturbing edges (Zügner and Günnemann, 2019; Xu et al., 2019; Zhang et al., 2022a; Wang et al., 2023). Some recent works (Sun et al., 2020; Wang et al., 2020; Zou et al., 2021; Ju et al., 2022a, b; Chen et al., 2022; Dai et al., 2020; Tao et al., 2021) study the node injection attack, which injects new nodes into a graph and connects them with some existing nodes. Chen et al. (Chen et al., 2022) prove that the node injection attack can theoretically cause more damage than the graph modification attack with less or equal modification budget. G-NIA model (Tao et al., 2021) sets a strong limitation that the attacker can only inject one node with one edge, and it achieves more than 90% successful rate in the single-target attack on Reddit and ogbn-products datasets. They demonstrate the strong potential of altering nodes’ existence, which is inspiring to our research. To the best of our knowledge, none of the adversarial attack methods considers node removal, since it is not common in attacking applications.
2.2. Graph Counterfactual Explanation
Graph counterfactual explanation aims to explain why a GNN model gives a particular result of a target node/edge/graph. Unlike the factual explanation that explains by associating the prediction with critical nodes or edges, the counterfactual explanation tries to find the minimum perturbation on the input graph that can change the prediction of the target. Some methods (Ying et al., 2019; Lucic et al., 2021; Bajaj et al., 2021; Luo et al., 2020) are purely based on edge removal; some methods utilize both node removal and edge removal, and the methods include optimizing mask matrices (Tan et al., 2022; Wu et al., 2021), predicting node influence (Rong et al., 2023), applying graph generation models (Ma et al., 2022; Yuan et al., 2020), or searching for an optimal neighbor graph (Yuan et al., 2021; Huang et al., 2023). As analyzed in Section 1 and shown in Table 1, our problem setting differs from existing works, so these methods are not directly applicable to our problem setting. We modify and adapt two widely used methods as supplementary baselines to this new problem. The first baseline is inspired by optimizing a mask matrix, which is a common method in graph counterfactual explanation (Lucic et al., 2021; Ying et al., 2019; Tan et al., 2022; Bajaj et al., 2021; Wu et al., 2021). They usually multiply the adjacency matrix with the mask matrix which indicates edge existence. During training, its elements are within $[0,1]$. During inference, elements below $0.5$ indicate edge removal. We adapt it to node removal by using a node mask vector. Our second baseline is inspired by a recent work, LARA (Rong et al., 2023), which trains a GCN model to predict node influence on the explanation target. The parameter size does not grow with the input graph size, so it is more scalable compared to previous methods. We adapt it as our second baseline.
3. Problem Definition
3.1. Notations
A graph $G=(V,E)$ consists of nodes $\boldsymbol{V}=\{v_{1},v_{2},...,v_{N}\}$ and edges $E=\{e_{ij}|j\in\mathcal{N}(i)\}$, where $\mathcal{N}(i)$ denotes the neighbor nodes of $v_{i}$ and $e_{ij}$ denotes the edge from $v_{i}$ to $v_{j}$. We use $\hat{\mathcal{N}}(i)$ to denote $\mathcal{N}(i)\cup\{v_{i}\}$. $\boldsymbol{A}$ denotes the adjacency matrix. Node $v_{i}$ is associated with feature vector $\boldsymbol{x_{i}}\in\mathbb{R}^{d}$, and a label $y_{i}\in\mathbb{R}$ if the node classification task is applicable. We denote the degree of $v_{i}$ as $d_{i}=|\mathcal{N}(i)|$. When we remove node $v_{r}\in V$, we also remove all edges connected with $v_{r}$ from graph $G$, and we denote the modified graph as $G_{-v_{r}}$. $g_{\theta}$ denotes a trained GNN model. We use $v_{r}$ to denote the target removing node and $\mathrm{F}_{g_{\theta}}(v_{r})$ to denote the influence of removing $v_{r}$.
Graph neural networks (GNNs) generally follow the message-passing framework (Gilmer et al., 2017). A GNN model consists of multiple graph convolutional layers. In a typical graph convolutional layer, a node updates its representation by aggregating its neighbor nodes’ representations:
(1) $\boldsymbol{h_{i}^{(l)}}=U_{l}(\boldsymbol{h_{i}^{(l-1)}},\mathrm{AGG}_{l}(\sum_{j\in\mathcal{\hat{N}}(i)}\mathrm{MSG}_{l}(\boldsymbol{h_{j}^{(l-1)}},\boldsymbol{h_{i}^{(l-1)}}))),$
where $\boldsymbol{h_{i}^{(l)}}$ denotes the representation of $v_{i}$ after the $l$-th layer ($l\in{1,2,\dots}$), and $\boldsymbol{h_{i}^{(0)}}$ is the input feature $\boldsymbol{x_{i}}$. $\mathrm{MSG}_{l}$ is the message function, $\mathrm{AGG}_{l}$ is the aggregation function, and $U_{l}$ is the update function.
3.2. Problem Definition
To evaluate the influence of node removal, we use GNN models as a surrogate to predict the change of nodes/edges caused by removing the target node. As shown in Figure 2, we measure the influence by the total variation distance between the original and updated predicted probability distribution, and we use the $\ell_{1}$-norm of the difference, which can equally capture the prediction change for every class.
Definition 1. (Node Influence in Node Classification Task) Given a node classification model $g_{\theta}$ trained on graph $G$, we denote its predicted class probability of node $v_{i}$ as $\boldsymbol{g_{\theta}(G)_{i}}\in\mathbb{R}^{c}$ ($c$ is the number of classes), the influence of node $v_{r}$ is calculated as:
(2) $\mathrm{F}_{g_{\theta}}(v_{r})=\sum_{i=1,i\neq r}^{N}||\boldsymbol{g_{\theta}(G)_{i}}-\boldsymbol{g_{\theta}(G_{-v_{r}})_{i}}||_{1}\,,$
Definition 2. (Node Influence in Link Prediction Task) Given a link prediction model $g_{\theta}$ trained on graph $G$, we denote its predicted probability of edge $e_{ij}$ as $f_{e}(g_{\theta}(G)_{e_{ij}})\in\mathbb{R}$, where $f_{e}(\cdot)$ is the optional layers that transform $g_{\theta}$’s represetations to the predicted edge probability. We use $D_{e}$ to denote the whole link prediction set, and $D_{r}$ to denote edges that connect to $v_{r}$. The influence of removing node $v_{r}$ is calculated as:
(3) $\mathrm{F}_{g_{\theta}}(v_{r})=\sum_{e_{ij}\in D_{e}-D_{r}}||f_{e}(g_{\theta}(G)_{e_{ij}})-f_{e}(g_{\theta}(G_{-v_{r}})_{e_{ij}})||_{1}\,,$
4. Methods
The task is to predict the defined influence score for every node on the graph. The biggest challenge is efficiency, and we want to avoid traversing all the nodes. We propose an intuitive, simple, yet effective method, NOde-Removal-based fAst GNN inference (NORA). In general, we approximate the node influence by analyzing the calculation formula and decomposing it into three parts, which correspond to three kinds of influence of the node removal. Then, we use gradient information and some heuristics to approximate it. Figure 3 illustrates the three kinds of influence. We mainly introduce our method in the node classification task, and after that, we will explain how to generalize it to the link prediction task.
Figure 3. Three kinds of influence of node removal: the disappearance of its node embedding; the change of its nearby nodes’ aggregation terms; and the spread-out influence to multi-hop neighbors.
4.1. Influence Score Calculation Decomposition
We cannot directly use the first-order derivatives to approximate node influence based on the definition in Equation 2, since there is a $\ell_{1}$-norm inside the summation. Intuitively, removing a node usually causes consistent change to the class of other nodes, e.g., raising/lowering the probability of some classes for all nodes. Thus, we approximate by moving the $\ell_{1}$-norm out of the summation. We denote the number of GNN layers as $L$, and $\boldsymbol{h_{i}^{(L)}}\in R^{c}$ is the predicted class probability of node $v_{i}$ ($c$ is the number of classes). We denote as $\boldsymbol{f_{r}}=\sum_{i=1,i\neq r}^{N}\boldsymbol{h_{i}^{(L)}}$ the sum of all nodes’ predictions except for node $v_{r}$, and we denote as $\boldsymbol{\delta f_{r}}$ the change of $\boldsymbol{f_{r}}$ caused by removing node $v_{r}$.
Lemma 0 ().
If removing $v_{r}$ consistently changes the class distributions of other nodes, its influence defined in Equation 2 is equal to:
(4) $||\sum_{i\neq r}(\boldsymbol{g_{\theta}(G)_{i}}-(\boldsymbol{g_{\theta}(G_{-v_{r}})_{i}})||_{1}=||\boldsymbol{\delta f_{r}}||_{1}=||\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}}\boldsymbol{\delta h_{i}^{(L)}}||_{1},$
where $\boldsymbol{\delta h_{i}^{(L)}}$ is the change of $\boldsymbol{h_{i}^{(L)}}$ caused by removing $v_{r}$. The formula above is strictly equal because $\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}}=\boldsymbol{1}$. We write it in this form because we want to keep a uniform form with later formulas. We can extend this form from the last layer to the previous layers. Here we analyze how to extend it from the $L$-th layer to the $(L-1)$-th layer. Equation 1 illustrates the framework of a message-passing GNN layer. We consider a typical parameterization of it:
(5) $\boldsymbol{h_{i}^{(l)}}=\sigma\big{(}\boldsymbol{W_{u}^{(l)}}(\boldsymbol{W_{s}^{(l)}}\boldsymbol{h_{i}^{(l-1)}}+\sum_{j\in\hat{\mathcal{N}}(i)}\alpha_{ji}^{(l)}\boldsymbol{W_{m}^{(l)}}\boldsymbol{h_{j}^{(l-1)}})\big{)},$
where $\sigma$ denotes the activation function, $\boldsymbol{W_{u}^{(l)}}$, $\boldsymbol{W_{s}^{(l)}}$, and $\boldsymbol{W_{m}^{(l)}}$ are model parameters. $\alpha_{ji}^{(l)}$ is the normalization term. The model parameters are fixed during inference. Therefore, we can approximate $\boldsymbol{\delta h_{i}^{(L)}}$ by the first-order derivatives as:
$\boldsymbol{\delta h_{i}^{(L)}}\approx-I(v_{r}\in\mathcal{N}(i))\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial h_{r}^{(L-1)}}}\boldsymbol{h_{r}^{(L-1)}}$
(6) $+\sum_{j\in\hat{\mathcal{N}}(i),j\neq r}(\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial\alpha_{ji}^{(L)}}}\delta\alpha_{ji}^{(L)}+\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial h_{j}^{(L-1)}}}\boldsymbol{\delta h_{j}^{(L-1)}}),$
where $I(.)$ is the indicator function. By incorporating the definition of $\boldsymbol{\delta f_{r}}$, we can derive the following formula.
Lemma 0 ().
We can approximate $\boldsymbol{\delta f_{r}}$ for the GNN model described in Equation 5 using the first-order derivatives as:
$\boldsymbol{\delta f_{r}}\approx-T_{1}+T_{2}+T_{3}=-\sum_{i\in\mathcal{N}(r)}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}}\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial h_{r}^{(L-1)}}}\boldsymbol{h_{r}^{(L-1)}}+\sum_{i\neq r}\sum_{j\in\hat{\mathcal{N}}(i),j\neq r}$
(7) $(\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}}\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial\alpha_{ji}^{(L)}}}\delta\alpha_{ji}^{(L)})+\sum_{i\neq r}\sum_{j\in\hat{\mathcal{N}}(r),j\neq r}(\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}}\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial h_{j}^{(L-1)}}}\boldsymbol{\delta h_{j}^{(L-1)}}).$
The calculation contains three terms. $T_{1}$ measures the disappearance of $v_{r}$’s latent representation as a message to its neighbor nodes; $T_{2}$ measures the change of its neighbors’ normalization term; $T_{3}$ measures the change of its neighbors’ latent representations. The three terms correspond to the three kinds of influence in Figure 3.
4.2. Approximation of Each Decomposed Term
$T_{1}$: Disappearance of the message to neighbor nodes. On the computation graph, $\boldsymbol{h_{r}^{(L-1)}}$ can connect to later layers either by $\boldsymbol{h_{r}^{(L)}}$ or by $\boldsymbol{h_{i}^{(L)}},i\in\mathcal{N}(r)$. Therefore, by applying the chain rule, the term $T_{1}$ is equal to:
(8) $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}}}\boldsymbol{h_{r}^{(L-1)}}-\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L)}}}\boldsymbol{\frac{\partial h_{r}^{(L)}}{\partial h_{r}^{(L-1)}}}\boldsymbol{h_{r}^{(L-1)}}.$
Although $\boldsymbol{h_{r}^{(L)}}$ is not related to $\boldsymbol{f_{r}}$ and $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L)}}}=\boldsymbol{0}$, we still write it in this way as a general form which can be later applied to previous layers, since $\boldsymbol{h_{r}^{(L-1)}}$, $\boldsymbol{h_{r}^{(L-2)}},etc.$ is related to $\boldsymbol{f_{r}}$. Equation 8 consists of two parts. The form of the first part is more convenient to handle, so we want to eliminate the second part. We do this by approximating the ratio of the second part to the first part. Here we make a rough assumption that every node is functionally and structurally equal, which means they have the same degree, the same representation, and the same gradient. We denote the gradient coming from a neighbor, $\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}}\boldsymbol{\frac{\partial h_{i}^{(L)}}{\partial h_{r}^{(L-1)}}}$, as $\boldsymbol{g}$, and the gradient coming from the higher-layer representation of a node itself, $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L)}}}\boldsymbol{\frac{\partial h_{r}^{(L)}}{\partial h_{r}^{(L-1)}}}$, as $\beta\boldsymbol{g}$. $\beta\in\mathbb{R}$ is usually larger than one due to self-loop and residual connection. Then, the ratio of the second part versus the first part in Equation 8 can be approximated as $\frac{\beta}{d_{r}+\beta}$, where $d_{r}$ is the degree of node $v_{r}$. Then we derive the following lemma.
Lemma 0 ().
If every node in the graph is structurally and functionally equal, we can approximate the term $T_{1}$ in Equation 7 as:
(9) $T_{1}\approx\frac{d_{r}}{d_{r}+\beta}\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}}}\boldsymbol{h_{r}^{(L-1)}}.$
For computation convenience, we approximate $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}}h_{r}^{(L-1)}}$ by a scalar. Based on experiments, we find that an effective way to approximate $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}}h_{r}^{(L-1)}}$ is calculating $||(\boldsymbol{f_{r}\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}})\circ h_{r}^{(L-1)}}||_{p}$, where $\circ$ is element-wise product and ||.$||.||_{p}$ is the $\ell_{p}$-norm. Here we multiply the original formula by $\boldsymbol{f_{r}}$ so that we increase the focus on the classes with high predicted probabilities. Here $\boldsymbol{f_{r}}\in\mathbb{R}^{c}$, $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}}}\in\mathbb{R}^{cxd}$ is the Jacobian matrix, and $\boldsymbol{h_{r}^{(L-1)}}\in\mathbb{R}^{d}$. p is a hyper-parameter, and in most cases, we set it to one.
$T_{2}$: Change of aggregation terms. There are two challenges in approximating the term $T_{2}$. First, calculating $\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}\frac{\partial h_{i}^{(L)}}{\partial\alpha_{ji}^{(L)}}}$ might consume too much space on large dense graphs with many edges. Second, the aggregation term differs significantly in different GNNs. For example, GCN (Kipf and Welling, 2016) and GraphSAGE (Hamilton et al., 2017) use node degree to compute it, while some models use the attention mechanism, such as GAT (Veličković et al., 2017). Therefore, it is difficult to derive a generally effective approximation.
After several attempts, we find that a widely effective way is to ignore $\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}\frac{\partial h_{i}^{(L)}}{\partial\alpha_{ji}^{(L)}}}$ and approximate $\delta\alpha_{ji}^{(L)}$ only by structure. We combine the aggregation term of GCN (Kipf and Welling, 2016) and GraphSAGE (Hamilton et al., 2017). The aggregation of GCN is $\alpha_{ji}^{(L)}=1/\sqrt{d_{i}d_{j}}$, and that of GraphSAGE (with mean aggregation) is $\alpha_{ji}^{(L)}=1/d_{i}$. If $v_{i}$ is not $v_{r}$’s neighbor, then we assume the aggregation term does not change. If $v_{i}$ is $v_{r}$’s neighbor, we approximate $\delta\alpha_{ji}^{(L)}$ by $\hat{\delta}\alpha_{ji}$ as:
$\hat{\delta}\alpha_{ji}=[k_{1}(\frac{1}{\sqrt{d_{i}-1}}-\frac{1}{\sqrt{d_{i}}})+(1-k_{1})(\frac{1}{d_{i}-1}-\frac{1}{d_{i}})]$
(10) $[k_{2}\frac{1}{\sqrt{d_{j}}}+k_{2}^{\prime}\frac{1}{d_{j}}+(1-k_{2}-k_{2}^{\prime})],$
where $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are hyper-parameters within [0,1]. There exist hyper-parameters that make $\hat{\delta}\alpha_{ji}$ satisfy GCN or GraphSAGE. Based on $\hat{\delta}\alpha_{ji}$, we approximate the second term as:
(11) $T_{2}\approx k_{3}\cdot\delta Topo_{r},\delta Topo_{r}=\sum_{i\in\mathcal{N}(r)}\sum_{j\in\mathcal{N}(i)}\hat{\delta}\alpha_{ji},$
where $k_{3}$ is a hyper-parameter.
$T_{3}$: Change of hidden representations of other nodes. When we analyze the term $T_{3}$, on the computation graph, $\boldsymbol{h_{j}^{(L-1)}}$ can reach $f_{r}$ by either $\boldsymbol{h_{r}^{(L)}}$ or $\boldsymbol{h_{i}^{(L)}},i\neq r$. Based on the chain rule, we can simplify $T_{3}$ in Equation 7 to Equation 12, which can be further transformed into Equation 13.
(12) $\sum_{j\neq r}(\boldsymbol{\frac{\partial f_{r}}{\partial h_{j}^{(L-1)}}}-\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L)}}}\boldsymbol{\frac{\partial h_{r}^{(L)}}{\partial h_{j}^{(L-1)}})}\boldsymbol{\delta h_{j}^{(L-1)}}$
(13) $=\sum_{j\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{j}^{(L-1)}}}\boldsymbol{\delta h_{j}^{(L-1)}}-\sum_{j\in\mathcal{N}(r)}\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L)}}\frac{\partial h_{r}^{(L)}}{\partial h_{j}^{(L-1)}}\delta h_{j}^{(L-1)}}.$
Although $\boldsymbol{h_{r}^{(L)}}$ is not related to $\boldsymbol{f_{r}}$ and $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(L)}}}=\boldsymbol{0}$, we still write it in this way as a general form which can be later applied to previous layers, since $\boldsymbol{h_{r}^{(L-1)}}$, $\boldsymbol{h_{r}^{(L-2)}},etc.$ is related to $\boldsymbol{f_{r}}$. Similar to the simplification process of $T_{1}$, here we also arrive at a formula with two parts. The form of the first part is more convenient to handle, and it takes the same form as Equation 4, so we want to eliminate the second part. We make the same rough assumption as approximating $T_{1}$ that every node is functionally and structurally equal. We denote the average node degree as $d$ and $\boldsymbol{\delta h_{j}^{(L-1)}}$ as $\boldsymbol{\delta h}$. Using the same notations of $\boldsymbol{g}$ and $\beta\boldsymbol{g}$ as approximating $T_{1}$, we approximate $\boldsymbol{\frac{\partial f_{r}}{\partial h_{j}^{(L-1)}}}$ as $(d+\beta)\boldsymbol{g}$, and we approximate the first part of Equation 13 as $(N-1)(d+\beta)\boldsymbol{g\delta h}$. We approximate the second part of Equation 13 as $d_{r}\boldsymbol{g\delta h}$. Then by rewriting the enumeration variable $j$ as $i$, we derive the following lemma.
Lemma 0 ().
If every node in the graph is structurally and functionally equal, we can approximate the term $T_{3}$ in Equation 7 as:
(14) $T_{3}\approx\big{(}\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L-1)}}}\boldsymbol{\delta h_{i}^{(L-1)}}\big{)}\big{(}1-\frac{d_{r}}{(N-1)(d+\beta)}\big{)}.$
4.3. Combined Derivation and Heuristics
By combining the approximations of three terms, we get:
(15) $\boldsymbol{\delta f_{r}}\approx\frac{d_{r}}{d_{r}+\beta}||\boldsymbol{(f_{r}\frac{\partial f_{r}}{\partial h_{r}^{(L-1)}})\circ h_{r}^{(L-1)}}||_{p}+k_{3}\cdot\delta Topo_{r}+T_{3}^{\prime},$
where $T_{3}^{\prime}$ is Equation 14. Here we remove the negative sign before $T_{1}$, because the original influence is based on the $\ell_{1}$-norm of the vector of the prediction change, but our approximation of the first and second term only results in a scalar. In reality, $-T_{1}$ might represent the decrease of the predicted probability of some classes that are related to the removed node, while $T_{2}$ might represent the increased predicted probability of other classes. To include both of their contributions instead of counteracting them, we directly add the approximated scalar for $T_{1}$ and $T_{2}$.
We successfully extend the original formula’s form from the L-th layer to (L-1)-th layer by Term $T_{3}$. By repeating this process, we can approximate $\boldsymbol{\delta f_{r}}$ by the gradient from every layer. We show the extension process in the appendix. By extending Equation 15 to all previous layers, we derive:
$\mathrm{F}_{g_{\theta}}(v_{r})\approx\sum_{i=0}^{L-1}(\hat{d}_{r}^{(L-1-i)}\hat{h}_{r}^{(i)})+k_{3}^{\prime}\cdot\delta Topo_{r},$
(16) $where\ \hat{d}_{r}=1-\frac{d_{r}}{(N-1)(d+\beta)},\quad\hat{h}_{r}^{(i)}=\frac{d_{r}}{d_{r}+\beta}||\boldsymbol{(f_{r}\frac{\partial f_{r}}{\partial h_{r}^{(i)}})\circ h_{r}^{(i)}}||_{p}.$
$\boldsymbol{h_{i}^{(0)}}$ is the input feature of node $v_{i}$. We aggregate $\delta Topo_{r}$ in all layers and change the hyper-parameter $k_{3}$ to $k_{3}^{\prime}$ as a result.
The way we derive Equation 16 closely corresponds to the three kinds of node influence in Figure 3. Approximation of term $T_{1}$ entails node embeddings and gradients, which corresponds to the disappearance of $v_{r}$’s embeddings. Approximation of term $T_{2}$ contributes to the structural influence $\delta Topo_{r}$, which corresponds to the change of aggregation terms. Approximation of term $T_{3}$ contributes to extending the formula from the last GNN layer to former layers, which enlarges the neighborhood size as it increases GNN layers and spreads out the influence.
Nonetheless, we cannot compute the approximation for all nodes simultaneously. We change $\boldsymbol{f_{r}}=\sum_{i=1,i\neq r}^{N}\boldsymbol{h_{i}^{(L)}}$ to the sum of all nodes’ predictions including $v_{r}$. In this way, all the nodes can share the same $f_{r}$. This will change the gradient $\boldsymbol{\frac{\partial f_{r}}{\partial h_{r}^{(i)}}}$, but this change might be similar to all nodes due to the following lemma.
Lemma 0 ().
If we remove the nonlinear activation function in the GNN layer in Equation 5, the gradient $\boldsymbol{\frac{\partial h_{r}^{(L)}}{\partial h_{r}^{(L-1)}}}$ can be calculated as:
(17) $\boldsymbol{\frac{\partial h_{r}^{(L)}}{\partial h_{r}^{(L-1)}}}=\boldsymbol{W_{u}^{(L)}(W_{s}^{(L)}}+\alpha_{rr}^{(L)}\boldsymbol{W_{m}^{(L)}}).$
As shown, the increased gradient could be highly relevant to the model parameters, so it might be similar among different nodes. Though it theoretically does not work on nonlinear GNNs, our experiments have shown a satisfactory performance. Now, we can approximate the influence score for all nodes simultaneously after only one forward computation and one backpropagation.
Approximation in the link prediction task is similar. We replace $\boldsymbol{f_{r}}$ in Equation 16 which is the sum of node predictions with the sum of edge predictions.
4.4. Complexity Analysis
We analyze the time and space complexity of the ground truth method (brute-force) and NORA. $N$ denotes node number, $M$ denotes edge number, $L$ denotes the number of GNN layers, and $h$ denotes the hidden size. In most cases, the adjacency matrix is sparsely stored, and in this situation, according to Blakely et al. (Blakely et al., 2021), the time complexity of one forward or backward propagation of a common message-passing GNN model is $O(LNh^{2}+LMh)$, and the space complexity is $O(M+Lh^{2}+LNh)$. We list the time and space complexities in Table 2. NORA costs significantly less time than the brute-force method, and basically the same space complexity, so it can be generalized to large real-world graphs.
Table 2. Complexity comparison.
Method Time Space
Brute-force $O(LN^{2}h^{2}+LNMh)$ $O(M+Lh^{2}+LNh)$
NORA $O(LNh^{2}+LMh)$ $O(M+Lh^{2}+LNh)$
5. Experiments
5.1. Baseline Adaption
There is no mature baseline for this new problem, so we adapt two methods from graph counterfactual explanation as baselines.
Node mask. Mask optimization is widely used in graph counterfactual explanation (Lucic et al., 2021; Ying et al., 2019; Tan et al., 2022; Bajaj et al., 2021; Wu et al., 2021). We use a mask vector $\boldsymbol{m}\in\mathbb{R}^{N}$ to indicate the existence of the $N$ nodes. Elements of $\boldsymbol{m}$ are limited in [0, 1]. In every GNN layer, we multiply node embeddings by $\boldsymbol{m}$ before the message passing. We fix the GNN’s parameters and only optimize the mask $\boldsymbol{m}$. Our optimization goal is to maximize the difference between the updated GNN’s prediction and its original prediction, calculated as the $\ell_{1}$-norm. After training, we evaluate the node influence as the distance between elements in $\boldsymbol{m}$ and $\boldsymbol{1}$. Following common designs, we use regularization terms. One regularization term drives the mask elements to zero, which guides elements in $\boldsymbol{m}$ to decrease instead of increase. The second regularization term drives the mask to one, without which the mask might become $\boldsymbol{0}$. Our loss function is:
(18) $Loss=-\sum_{i=1}^{N}||\boldsymbol{g_{\theta}(V,E)_{i}}-\boldsymbol{g_{\theta}(V,E,m)_{i}}||_{1}+\alpha||\boldsymbol{m}||_{1}+\beta||\boldsymbol{1}-\boldsymbol{M}||_{1}.$
Prediction model. A recent work, LARA (Rong et al., 2023), greatly improves the counterfactual explanation methods’ scalability by applying a GNN model to predict the node/edge influence on the explanation target, so the parameter size is agnostic with the graph size. Inspired by LARA, we train a GCN model to generate a source embedding, $\boldsymbol{p_{i}}$, and a target embedding, $\boldsymbol{t_{i}}$ for every node $v_{i}$. We predict the influence of node $v_{i}$ on node $v_{j}$ by $\boldsymbol{p_{i}}\cdot\boldsymbol{t_{j}}$, where $\cdot$ is the dot product. We predict the influence of a node as the sum of the predicted influence of its outgoing edges as $\mathrm{F}_{g_{\theta}}(v_{r})\approx\sum_{i\in N(r)}\boldsymbol{p_{r}}\cdot\boldsymbol{t_{i}}$.
Besides, we also try to directly predict the node influence score by a GNN model, which is a node regression task. In the following tables, “Predict-E” is the first way, and “Predict-N” is directly predicting node influence. We have tried different GNN models to do the prediction, including GCN and GAT, while GCN performs the best. It might be because we use a few-shot setting. With a very limited number of labels to train the model (7%), complex GNN structures like GAT might not be well-trained.
Table 3. Dataset statistics.
Dataset #Nodes #Edges #Features #Classes Homo/Hetero
Cora 2,708 5,429 1,433 7 homogeneous
CiteSeer 3,327 4,732 3,703 6 homogeneous
PubMed 19,717 44,338 500 3 homogeneous
ogbn-arxiv 169,343 1,166,243 128 40 homogeneous
P50 5,435 1,593,721 one-hot 2 heterogeneous
P_20_50 12,103 1,976,985 one-hot 2 heterogeneous
Table 4. Pearson correlation coefficient between real influence scores and approximated scores on citation datasets.
Node classification Link prediction
GNN Model Method Cora CiteSeer PubMed ogbn-arxiv Cora CiteSeer PubMed ogbn-arxiv
GCN Predict-N 0.737 0.749 0.896 0.873 0.811 0.777 0.901 0.655
Predict-E 0.788 0.703 0.823 0.800 0.859 0.735 0.901 0.842
Node mask 0.880 0.864 0.900 0.847 0.942 0.871 0.922 0.908
NORA-$T_{1}$ 0.876 0.831 0.847 0.899 0.850 0.848 0.851 0.945
NORA-$T_{2}$ 0.869 0.829 0.927 0.952 0.946 0.911 0.947 0.977
NORA 0.903 0.901 0.927 0.956 0.967 0.926 0.949 0.977
GraphSAGE Predict-N 0.712 0.709 0.808 0.856 0.693 0.595 0.877 0.52
Predict-E 0.775 0.794 0.792 0.833 0.930 0.891 0.923 0.835
Node mask 0.825 0.892 0.896 0.878 0.816 0.305 0.948 0.734
NORA-$T_{1}$ 0.829 0.819 0.816 0.943 0.944 0.903 0.813 0.898
NORA-$T_{2}$ 0.859 0.838 0.831 0.956 0.923 0.842 0.933 0.927
NORA 0.896 0.889 0.860 0.957 0.978 0.934 0.971 0.936
GAT/DrGAT Predict-N 0.690 0.722 0.867 0.685 0.734 0.726 0.844 0.526
Predict-E 0.842 0.754 0.764 0.777 0.918 0.857 0.910 0.845
Node mask 0.878 0.834 0.836 0.783 0.952 0.860 0.906 0.617
NORA-$T_{1}$ 0.916 0.828 0.829 0.147 0.958 0.815 0.877 0.799
NORA-$T_{2}$ 0.891 0.886 0.907 0.909 0.930 0.927 0.862 0.828
NORA 0.927 0.904 0.910 0.909 0.982 0.933 0.951 0.884
GCNII Predict-N 0.729 0.739 0.824 0.873 0.794 0.777 0.882 0.718
Predict-E 0.740 0.769 0.816 0.765 0.912 0.809 0.914 0.822
Node mask 0.860 0.881 0.898 0.827 0.946 0.867 0.935 0.733
NORA-$T_{1}$ 0.702 0.884 0.828 0.910 0.931 0.804 0.875 0.940
NORA-$T_{2}$ 0.811 0.908 0.847 0.953 0.962 0.903 0.948 0.987
NORA 0.874 0.919 0.874 0.957 0.969 0.916 0.957 0.987
Table 5. Pearson correlation coefficient between real influence scores and approximated scores on Twitter datasets.
Method P50 node P_20_50 node P50 link P_20_50 link
Predict-N 0.405 0.119 0.526 0.724
Predict-E 0.738 0.727 0.791 0.806
Node mask 0.971 0.652 0.968 0.942
NORA-$T_{1}$ 0.951 0.751 0.910 0.903
NORA-$T_{2}$ 0.625 0.764 0.684 0.813
NORA 0.953 0.849 0.912 0.914
5.2. Experiment Settings
Datasets. We conduct extensive experiments on six datasets. They include four widely used benchmark citation datasets (Cora, CiteSeer, and PubMed (Sen et al., 2008), and ogbn-arxiv (Hu et al., 2020)) and two Twitter datasets (P50 and P_20_50 (Xiao et al., 2020)). The four citation networks are homogeneous undirected graphs. A node is a paper, and an undirected edge represents a citation. The original task is to predict the research field of each paper. We add a link prediction task, and we use the dot product of two nodes’ representations plus a sigmoid function to do the prediction. The two Twitter datasets are heterogeneous directed graphs. Nodes are users and directed edges represent one of five Twitter interactions or their counterparts (e.g., be followed): follow, retweet, like, reply, and mention. It has both node classification (predicting the political leaning) and link prediction, so we use their original tasks. Table 3 lists the dataset statistics. An issue is that the trained GNN model is biased to the training-set nodes/edges. To fairly evaluate node influence, we run each experiment 5 times and cycle the data split of nodes/edges by 20% per time, giving every node/edge an equal chance to show up in training, validation, or test sets. We take the mean of the 5 results.
GNN Models. We select six representative GNN models. On the four citation datasets, we use GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2017), and GCNII (Chen et al., 2020). As the ogbn-arxiv dataset is a heated OGB benchmark, we use the SOTA model at the time we started this project, DrGAT (Zou et al., 2019), to replace GAT in the node classification task. DrGAT is an improved variant of GAT, which is equipped with a dimensional reweighting mechanism. Since the two Twitter datasets are heterogeneous, the above models can no longer be directly applied, so we use TIMME model which is proposed in the same paper as the Twitter datasets (Xiao et al., 2020). TIMME tackles three challenges on the Twitter datasets: sparse feature, sparse label, and heterogeneity. We use the hyper-parameters for DrGAT and TIMME models in their GitHub repository, since their hyper-parameters have been carefully selected. We tune the hyper-parameters for GCN, GraphSAGE, GAT, and GCNII. We also tune the hyper-parameters of each approximation method for each dataset and model.
Evaluation. Label generation by the brute-force method is time-consuming, so we evaluate the methods’ performance in the few-shot setting, which is more suitable for real-world applications. The methods only have access to 10% of real influence scores, and the other 90% are for testing. The “node mask” method and NORA use the 10% as the validation set to tune hyper-parameters. The “prediction” method uses 7% to train and 3% for validation. The evaluation metric is the Pearson correlation coefficient between the real influence and the approximated/predicted influence.
5.3. Performance Comparison
Table 4 shows the results on the four citation datasets, and Table 5 shows the results on the two Twitter datasets. NORA outperforms the baseline methods in most cases, demonstrating its effectiveness. Among the baseline methods, the “node mask” method performs the best. It is more useful in its original design, which is to analyze the influence of a few nodes or edges. When we need to model the influence of all nodes, different nodes/edges are dominantly influenced by different nodes, so it is more difficult to optimize. The “prediction” method is greatly limited by label usage. It requires additional ground truth node influence to train, but since we only have a very small training set of 7% nodes, its potential is limited. We analyze the results of enlarging the training set in the appendix. If we increase the label usage, its performance could be better, but NORA still significantly outperforms it.
Besides, we evaluate two variations of NORA as ablation studies. If we only consider the first term $T_{1}$ and the third term $T_{3}$ in the approximation (represented as NORA-$T_{1}$ in the tables), we will only consider the task-specific influence of the embeddings but ignore the structural influence, which equals setting $k_{3}^{\prime}$ to zero in Equation 16. If we only consider the second term $T_{2}$ and the third term $T_{3}$ (represented as NORA-$T_{2}$), we will only consider the structural influence, which equals to setting $k_{3}^{\prime}$ to infinity in Equation 16. NORA outperforms both variants, demonstrating the benefit of ensembling the task-specific influence and structural influence. In some situations, only considering one of them could already have a good performance that is close to NORA. It indicates that sometimes the influence is dominated by the task-specific influence or the structural influence. Ensembling both of them provides us with the opportunity to balance between them using the hyper-parameter $k_{3}^{\prime}$.
Here we intuitively analyze the approximation errors of NORA. The inaccuracy comes from these sources: (1) we only use the first-order derivative for approximation; (2) we use the aggregation term of GCN and GraphSAGE to approximate the aggregation term change. It is inaccurate for more complex aggregation methods like the attention mechanism in GAT; (3) Equation 9 and Equation 14 are derived from the rough assumption that every node is functionally and structurally equal, which is not the reality. If the nodes are more diverse, the approximation could be less accurate. There are also other sources of approximation error, such as using the sum of all output predictions instead of $\boldsymbol{f_{r}}$, etc. Our analysis could only provide a reference, but the factors are indeed very complex. It is difficult to predict the performance of NORA given a new dataset or a new GNN model. NORA contains several inaccurate and intuitive approximations. Nonetheless, its advantage is very high efficiency, and the experiment results on six GNN models and six datasets have already demonstrated its effectiveness.
5.4. Case Study
The node classification task on the ogbn-arxiv dataset is to classify each node (paper) into one of forty CS fields defined by the arXiv category (https://arxiv.org/category_taxonomy). The top-10 influential nodes in the dataset, evaluated by DrGAT model, are listed as below: Adam (Kingma and Ba, 2014), ResNet (He et al., 2016), VGGNet (Simonyan and Zisserman, 2014), an important improvement to Skip-Gram (Mikolov et al., 2013b), ImageNet (Russakovsky et al., 2015), the paper that proposed word embedding (Mikolov et al., 2013a), GoogLeNet (Szegedy et al., 2015), the paper that proposed the batch normalization (Ioffe and Szegedy, 2015), the Caffe framework (Jia et al., 2014), and Faster R-CNN (Ren et al., 2015). They are all well-known papers that revolutionized the related fields. It is reasonable that removing them would result in the change of predicted categories of related papers.
5.5. Time Consumption
Table 6. Time of calculating the ground truth.
Node classification Link prediction
Dataset GCN SAGE (Dr)GAT GCNII GCN SAGE GAT GCNII
Cora 41s 23s 42s 47s 35s 25s 61s 48s
CiteSeer 67s 39s 46s 70s 44s 35s 73s 100s
PubMed 403s 382s $\approx$15min $\approx$15min 391s 425s $\approx$12min $\approx$17min
ogbn-arxiv $\approx$9h $\approx$9h $\approx$41h $\approx$10h $\approx$2.5h $\approx$3h $\approx$13h $\approx$10h
P50 TIMME model: 14min TIMME model: $\approx$1.5h
P_20_50 TIMME model: $\approx$41min TIMME model: $\approx$4h
To provide a reference of time consumption, we list the time cost of the brute-force method to calculate the real node influence scores in Table 6. The time cost is positively related to the graph size, and the ogbn-arxiv dataset containing 169,343 nodes costs the longest time. It takes about 41 hours for the DrGAT model on the ogbn-arxiv dataset. In comparison, NORA and the two “prediction” baseline methods only need less than one minute. The “node mask” baseline method takes less than six minutes. Since we use 10% of labels as the training and/or validation sets, the total time cost is dominated by generating the ground truth on large graphs.
5.6. Stability of the Proposed Influence Score
We want to examine whether this new definition of node influence is stable. We evaluate the stability of the real node influence across different GNNs and different hyper-parameters. We choose an important hyper-parameter, hidden size. We conduct experiments on the node classification task on the four citation datasets. we use GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2017) in this experiment. As in previous experiments, we replace the GAT model with the DrGAT model on ogbn-arxiv. For each model, we use three different hidden sizes: 128, 256, and 512, except for DrGAT on ogbn-arxiv, which only uses 128 and 256 due to memory limitations. For each model and each dataset, we calculate the Pearson correlation coefficient of the influence scores of every pair of different hidden sizes, and we report the mean of those correlation coefficients. These results are in the left three columns in Table 7. When measuring the cross-GNN stability, for each hidden size and each dataset, we calculate the Pearson correlation coefficient of the influence scores of every pair of two different GNNs, and we report the mean of them in the last column (“Inter-model”) in Table 7.
Table 7. Stability results.
Dataset GCN GraphSAGE GAT/DrGAT Inter-model
Cora 0.9956 0.9857 0.9393 0.8765
CiteSeer 0.9968 0.9931 0.9585 0.8167
PubMed 0.9970 0.9963 0.9451 0.8372
ogbn-arxiv 0.9984 0.9979 0.9914 0.9557
The node influence scores generated by the same GNN with different hidden sizes are quite similar, which demonstrates the stability of the proposed node influence score. The influence scores generated by different GNNs are less similar. In the ideal case, if the GNNs could accurately capture the underlying information spreading patterns of the original graph, then their generated node influence scores should be the same. However, they can not achieve 100% accuracy. How much the calculated node influence could reveal the actual node influence depends on the accuracy of the GNN model as our surrogate to capture the real information spreading patterns.
6. Conclusion
We provide a new perspective of evaluating node influence: the task-specific node influence on GNN model’s prediction based on node removal. We use graph neural network (GNN) models as a surrogate to learn the underlying message propagation patterns on a graph. After training a GNN model, we remove a node, apply the trained GNN model, and use the output change to measure the influence of the removed node. To overcome the low efficiency of the brute-force method (ground truth), we analyze how GNN’s prediction changes when a node is removed, decompose it into three terms, and approximate them with gradients and heuristics. The proposed method NORA can efficiently approximate node influence for all nodes after only one forward propagation and one backpropagation. We conduct extensive experiments on six networks and demonstrate NORA’s effectiveness and efficiency. No matter how we evaluate node influence, we can never touch the real node influence but can only model it, so the modeling perspective is very important. This paper provides a novel perspective of evaluating node influence and offers an intuitive, simple, yet effective solution. Future works are required to better understand node influence and improve the approximation performance.
Acknowledgements.
This work was partially supported by NSF 2211557, NSF 1937599, NSF 2119643, NSF 2303037, NSF 20232551, NASA, SRC JUMP 2.0 Center, Cisco research grant, Picsart Gifts, and Snapchat Gifts.
Appendix A Supplementary Methods
In Section 4.3, we combine the approximation of the three terms to form Equation 15, then we extend it to fronter layers and acquire Equation 16. Here we explain how to extend the formula to previous layers. In the $T_{3}^{\prime}$ in Equation 15, we have $\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L-1)}}\delta h_{i}^{(L-1)}}$. Its form is very similar to the form of $\delta f_{r}$:
(19) $\boldsymbol{\delta f_{r}}=\sum_{i\neq r}\boldsymbol{\delta h_{i}^{(L)}}=\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}\delta h_{i}^{(L)}}.$
We can use the same method of approximating $\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(L)}}\delta h_{i}^{(L)}}$ to approximate $T_{3}^{\prime}$. After this approximation, we extend the formula to the (L-2)-th layer. By repeating this approximation method layer by layer, we can approximate previous layers similarly and extend the formula to previous layers. When we reach the first GNN layer (the layer after the input), we get:
$\mathrm{F}_{g_{\theta}}(v_{r})\approx\sum_{i=0}^{L-1}\big{(}\hat{d}_{r}^{(L-1-i)}(\hat{h}_{r}^{(i)}+k_{3}\cdot\delta Topo_{r})\big{)}+\hat{d}_{r}^{L}\cdot(\sum_{i\neq r}\boldsymbol{\frac{\partial f_{r}}{\partial h_{i}^{(0)}}\delta h_{i}^{(0)}}),$
(20) $where\ \hat{d}_{r}=1-\frac{d_{r}}{(N-1)(d+\beta)},\quad\hat{h}_{r}^{(i)}=\frac{d_{r}}{d_{r}+\beta}||\boldsymbol{(f_{r}\frac{\partial f_{r}}{\partial h_{r}^{(i)}})}\circ\boldsymbol{h_{r}^{(i)}}||_{p}.$
In the formula, $\boldsymbol{h_{i}^{(0)}}$ is the input feature of $v_{i}$. It does not change when another node is removed, so $\boldsymbol{\delta h_{i}^{(0)}}=\boldsymbol{0}$. Besides, since $\delta Topo_{r}$ is the same in every layer and is only determined by the graph structure, we extract it from the summation, and we re-assign its weight to be $k_{3}^{\prime}$. In this way, we can get the final formula of approximating node influence as Equation 16.
Appendix B Supplementary Experiments
B.1. Enlarging Label Usage of Baseline
As we analyzed previously, the baseline method of predicting node influence by GCN is greatly limited by label scarcity due to our few-shot setting. We explore the influence of enlarging label usage on its performance. We conduct experiments on the (Dr)GAT model on the ogbn-arxiv dataset whose label generation takes the longest time. We use the DrGAT (Zou et al., 2019) model for node classification and the GAT (Veličković et al., 2017) model for link prediction. We keep the ratio of training: validation as 7:3, and we increase the label usage ratio from 10% to 20%, 30%, and 40%. We list the performance in Table 8. In most cases, the model performance increases when there are more labels until the label number has already been sufficient (e.g., 40%). With the increased performance, the label generation time also increases a lot. In contrast, NORA does not require training and only acquires a small validation set to tune hyper-parameters, so it is more scalable on large graphs.
Table 8. Performance with more labels.
DrGAT (node classification) GAT (link prediction)
Method 10% 20% 30% 40% 10% 20% 30% 40%
Predict-N 0.685 0.698 0.718 0.780 0.526 0.535 0.557 0.578
Predict-E 0.777 0.802 0.812 0.795 0.845 0.777 0.828 0.693
B.2. Structural Patterns on Twitter Datasets
Figure 4. Relationship between node influence and degree.
Since the two Twitter datasets are directed heterogeneous graphs (the four citation datasets are undirected homogeneous graphs), it is worth studying the relationship between structural patterns and node influence on the two Twitter datasets. We first study the relationship between node influence and in-degree or out-degree. For each node, we calculate the mean real influence scores of the two tasks (node classification and link prediction), and we divide the nodes into 100 groups according to their influence level. Then we calculate each group’s average in-degree, out-degree, and the ratio of in-degree versus out-degree. We show the results in Figure 4. Low-influential users have small in-degrees and out-degrees. They might rarely use Twitter. As influence grows, in-degree significantly grows (the light blue area in Figure 4), while out-degree does not significantly grow from medium-influential to high-influential users (the deep blue area in Figure 4). It is reasonable since high-influential people get a lot more attention from others than medium-influential people, but high-influential people don’t necessarily pay a lot more attention to others. On the P50 dataset, as the influence grows, the ratio of in-degree versus out-degree also significantly grows. However, the ratio of in-degree versus out-degree of the P_20_50 dataset is abnormal. The ratio is high in the low-influential groups. It might be due to the randomness when both in-degree and out-degree are very small.
Figure 5. The relationship between node influence and edge type. “Ratio” means the mean node degree in a certain influence level divided by the total number of edges. Here, the degree and number of edges are separately computed for each edge type.
We also analyze the relationships between node influence and edge type. On the two Twitter datasets, there are 10 types of directional edges: reply, follow, retweet, mention, like, replied by others, followed by others, retweeted by others, mentioned by others, and liked by others. We divide all the nodes into 20 groups according to their influence levels based on the real influence scores. We also use the mean influence scores of the two tasks (node classification task and link prediction task). For every edge type, we calculate the mean degree of nodes within each influence level. Here, when we calculate the node degree for one edge type, we only consider edges of that type. Since different edge types have significantly different numbers of edges, we normalize the node degree by the total number of edges of the corresponding edge type. Since we have separated the “follow” relation with “followed”, “reply” with “replied”, etc., we no longer separately calculate in-degree and out-degree here.
We plot the results in Figure 5. On the P50 dataset, the node degree of “replied” and “followed” have the strongest positive correlation with the node influence. This observation coincides with TIMME paper (Xiao et al., 2020)’s observation that “follow/followed” and “reply/replied” are the two most important relations to predict a user’s political leaning. On the P_20_50 dataset, “follow” replace “followed” as the strongest indicator. This might be because the P_20_50 dataset is less politics-centered than the P50 dataset. The P50 dataset contains politicians and users who follow or are followed by no less than 50 politicians. The most influential users are probably some politicians, and the number of their followers could be a strong indicator of their influence. In contrast, the P_20_50 dataset contains politicians and users who follow or are followed by no less than 20 and less than 50 politicians. The most influential users for GNN’s prediction are probably those who follow many politicians and their supporting groups. People usually follow people with the same ideology, so the “follow” relationship can be very influential for the TIMME model’s prediction, whose node classification target is to predict users’ political leanings. Once we remove an account that follows many other people, it might not significantly change other people’s political leanings in reality, but the GNN model’s predictions might have a big change.
B.3. Uneven Distribution of Influence
A real-world application of evaluating task-specific node influence is to use a small number of nodes to trigger a big impact, such as viral advertising (Domingos and Richardson, 2001; Nguyen et al., 2017; Leskovec et al., 2007a), online news dissemination (Leskovec et al., 2007b; Gruhl et al., 2004), finding the bottlenecks in an infrastructure network to improve its robustness (Li and Liu, 2021; Crucitti et al., 2004), etc. In order to achieve these goals, an important feature is that a small number of nodes have a large influence, compared to most nodes having a small influence. Here we analyze the divergence of the proposed node influence. We calculate the ratio of the summed influence of the top kinfluential nodes compared to the sum of all nodes’ influence. We calculate the mean results of the real influence scores generated by all GNN models we have used on each dataset. Table 9 shows the results. A small portion of top influential nodes have a large influence. Node influence on the two Twitter datasets is more unevenly distributed than the four citation datasets. Only 10% of people contribute to more than 75% of influence on the node classification task on the Twitter datasets. It is reasonable as a small portion of people on social media attract much more attention than most people. For most datasets, the influence on the node classification task is more unevenly distributed than on the link prediction task. It might be because the prediction of an edge is based on two node embeddings, but the prediction of a node is only based on its own embedding, so node classification might be more sensitive to single-node-removal perturbation than link prediction.
Table 9. The ratio (%) of top kinfluential nodes’ summed influence divided by the summed influence of all nodes. “Node” represents the node classification task; “Edge” represents the link prediction task.
kCora CiteSeer PubMed ogbn-arxiv P50 P_20_50
Node 1% 14.78 8.64 15.32 11.08 29.42 38.85
3% 24.95 17.15 30.38 17.84 45.93 60.71
10% 45.14 35.90 57.20 31.78 77.26 78.39
Edge 1% 12.01 8.24 10.68 19.78 21.80 24.69
3% 19.58 15.97 21.93 29.75 34.84 42.13
10% 34.46 32.60 45.48 46.69 60.20 63.97
B.4. Experiment Details
Data split. For the node classification task, we use the original data split ratio for ogbn-arxiv, P_50, and P_20_50. For Cora, CiteSeer, and PubMed, in the original data split, the majority of nodes are not in any of the training set, validation set, or test set in the original data split, so we change the data split ratio to train:valid:test = 5:3:2 to cover all nodes.
For the link prediction task, we use the original data split ratio for P50 and p_20_50, which is train:valid:test = 8:1:1, and it randomly samples three times negative edges as positive edges. Since there is no original link prediction task on Cora, CiteSeer, PubMed, and ogbn-arxiv datasets, we implement the link prediction task ourselves. We use the same data split ratio of 8:1:1. We only randomly sample the same number of negative edges as positive edges, because we find that too many negative edges on these four datasets would result in model collapse, which simply predicts “no edge”. It might be because we only implement a simple model, which calculates the dot product of two node embeddings plus a sigmoid function. During training, the model can access the training-set edges, and it needs to predict both the training-set edges and training-set negative edges. During evaluation, the model can access the training-set edges, and it needs to predict the edges and negative edges in the validation set or test set.
Hyper-parameter tuning. We tune the hyper-parameters for every GNN model and every approximation/prediction method. We use six GNN models: GCN, GraphSAGE, GAT, DrGAT, GCNII, and TIMME. We tune hyper-parameters for GCN, GraphSAGE, GAT, and GCNII. For the DrGAT model and the TIMME model, we use their original hyper-parameters provided by their GitHub repositories, since they have already been carefully tuned by their original papers. Please refer to DrGAT’s repository  [^1] https://github.com/anonymousaabc/DRGCN and TIMME’s repository  [^2] https://github.com/PatriciaXiao/TIMME for more details. We use two-layer GCN, GraphSAGE, GAT, and GCNII when applied to Cora, CiteSeer, and PubMed, and three layers when applied to ogbn-arxiv. On the four citation datasets, the last GNN layer directly provides the output predictions for the node classification task. For the link prediction task, we use the same hidden size in the last layer’s output as in hidden layers, and we use the dot product between two nodes’ outputs plus a sigmoid function to generate the prediction. The TIMME model contains two GNN layers followed by a node classification MLP or a link prediction module. During hyper-parameter tuning, we mainly tune the hidden size, dropout, learning rate, and weight decay, and we tune the hyper-parameters for every GNN model on every dataset. We store the hyper-parameters in our GitHub repository (https://github.com/weikai-li/NORA.git). Following our notes and running the default codes will automatically use the selected hyper-parameters.
We also tune the hyper-parameters for NORA and every baseline method, and we tune them for each dataset and each GNN model. NORA has five hyper-parameters: $k_{1}$, $k_{2}$, $k_{2}^{\prime}$, $k_{3}^{\prime}$, and $\beta$. Among them, $\beta$ is related to approximating term $T_{1}$ and term $T_{3}$; $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are related to approximating term $T_{2}$; $k_{3}^{\prime}$ is used to adjust the weight between the two components in Equation 16. For convenience, we normalize the two components before multiplying $k_{3}^{\prime}$. Based on our experience, we can separately tune the hyper-parameters for the two components, and the best hyper-parameter setting is usually the combination of the best hyper-parameter setting for the two components. $\beta$ is usually within [1, 20]; $k_{1}$, $k_{2}$, and $k_{2}^{\prime}$ are within [0, 1]; $k_{3}^{\prime}$ is usually within [0.5, 5] and not far from one, since we normalize the two components.
The “node mask” baseline method has four hyper-parameters to tune: the learning rate, the two weights $\alpha$ and $\beta$ of its two regularization terms, and the number of training epochs. We usually optimize the mask for about 100 to 300 epochs, and we use the mask vector which achieves the lowest validation loss (not including the regularization loss).
The “prediction” baseline method which trains a GCN model to predict node influence has six hyper-parameters to tune: the learning rate, weight decay, number of training epochs, hidden size, number of GCN layers, and the dropout. We usually train the model for 100 to 300 epochs, and we use the model that has the lowest validation loss during training. We use the MSE loss function for this regression task.
We store all the hyper-parameters in our GitHub repository (https://github.com/weikai-li/NORA.git), and we provide scripts for every approximation method for every GNN model and every dataset. With the scripts containing the hyper-parameters, it is easy to reproduce our results.